OpenStudy (anonymous):

Estimate the product of 7.8 and 11.4 by rounding.

OpenStudy (chaise):

7.8 rounds to 8 and 11.4 rounds to 11. The product of two numbers is their multiplication 8 * 11 = 88 An esitmation of 7.8 * 11.4 =~ 88 by roudning.

OpenStudy (anonymous):

ok so 7.8 is 8 and 11.4 is 11 the product means multiply therefore it would be 8*11=88

OpenStudy (chaise):

Correct :)

OpenStudy (anonymous):

Yep :) thanks for the medal :D